# Adaptive Randomized Smoothing: Certifying Multi-Step Defences against Adversarial Examples

### v1

@@ Abstract
We propose Adaptive Randomized Smoothing (ARS) to certify the predictions of our test-time adaptive models against adversarial examples. ARS extends the analysis of randomized smoothing using f-Differential Privacy to certify the adaptive composition of multiple steps. For the first time, our theory covers the sound adaptive composition of general and high-dimensional functions of noisy input. We instantiate ARS on deep image classification to certify predictions against adversarial examples of bounded $L_{\infty}$ norm. In the $L_{\infty}$ threat model, our flexibility enables adaptation through high-dimensional input-dependent masking. We design adaptivity benchmarks, based on CIFAR-10 and CelebA, and show that ARS improves accuracy by $2$ to $5\%$ points. On ImageNet, ARS improves accuracy by $1$ to $3\%$ points over standard RS without adaptivity.
@@ 1 Introduction
Despite impressive accuracy, deep learning models still show a worrying susceptibility to adversarial attacks. Such attacks have been shown for a large number of tasks and models (Costa et al., 2023; Chakraborty et al., 2018), including areas where security and safety are critical such as fraud detection (Pumsirirat and Liu, 2018) or self-driving (Cao et al., 2021).
@@
We evaluate our adaptive randomized smoothing method in three settings (§4). For image classification, we first design a challenging benchmark based on CIFAR-10, and we show that ARS can improve accuracy by up to $9.9\%$. For spatially-localized face attribute classification on the CelebA dataset, we show that ARS improves accuracy by up to $10\%$. For large-scale image classification on ImageNet, ARS improves the accuracy by $1\%$. At radius $0.005$, where RS has certified accuracy of 00 while ARS still has a certified accuracy of $15\%$.
@@ 2.1 Related Work: Adversarial Robustness, Randomized Smoothing, and Differential Privacy
While sound, RS is static during testing even though attacks may adapt. Recent work aims to make RS adapt at test time (Súkeník et al., 2021; Alfarra et al., 2022a; Hong et al., 2022). While pioneering, these works are restricted in either their soundness or their degree of adaptation and resulting improvement. Súkeník et al. (2021) soundly adapt the variance for RS by the distance between test and train inputs. However, this only provides minimal adaptivity, with only minor improvement to certification. Alfarra et al. (2022a) adapt the variance for RS to each test input, but the analysis is not end-to-end, and hence not sound (Súkeník et al., 2021). UniCR (Hong et al., 2022) adapts the noise distribution for RS, primarily to the data distribution during training, and optionally to input during testing. The train-time adaptation is sound, but the test-time adaptation is not due to the same issue raised by Súkeník et al. (2021). We propose ARS to advance certified test-time adaptivity: our approach is sound and high-dimensional to flexibly adapt the computation of later steps conditioned on earlier steps by differential privacy theory.
@@
$\boldsymbol{f}$-DP (Dong et al., 2019) is a notion of privacy that extends $(\epsilon,\delta)$-DP, and defines privacy as a bound on the power of hypothesis tests. Appendix A provides more details on $f$-DP. The main result we leverage is Theorem 2.7 of Dong et al. (2022) that for a Gaussian mechanism $\mathcal{M}(X)=\theta(X)+z$, $z\sim\mathcal{N}(0,\frac{r^{2}}{\mu^{2}})$, such that for any neighbouring $X,X^{\prime}$, $\theta(X)-\theta(X^{\prime})\in B_{2}(r)$ (i.e., the $L_{2}$ sensitivity of $\theta$ is $r$), we have that $\mathcal{M}$ is $G_{\mu}$-DP with:
$G_{\mu}(\alpha)=\Phi\Big{(}\Phi^{-1}(1-\alpha)-\mu\Big{)}.$ (2)
We leverage two key properties of $f$-DP. First, $f$-DP is resilient to post-processing. That is, if mechanism $\mathcal{M}$ is $f$-DP, $\textrm{proc}\circ\mathcal{M}$ is also $f$-DP. Second, $f$-DP is closed under adaptive composition. We refer to §3 in Dong et al. (2022) for the precise definition and use their Corollary 3.3: the adaptive composition of two Gaussian mechanisms $G_{\mu_{1}}$-DP and $G_{\mu_{2}}$-DP is itself $G_{\mu}$-DP, with:
$\mu=\sqrt{\mu_{1}^{2}+\mu_{2}^{2}}$ (3)
@@ 2.4 ARS against $L_{\infty}$-Bounded Adversaries
Important remarks. 1. $w(.)$ is a masking function, adaptively reducing (if $w_{i}(m_{1})\ll 1$) the value of $X_{i}$ and thereby the attack surface of an $L_{\infty}$ attacker. This reduces the effective dimension of the input to $\mathcal{M}_{2}$. 2. Reducing the dimension enables a reduction in the noise variance in $\mathcal{M}_{2}$, at fixed privacy guarantee $G_{\mu_{2}}$. The variance reduction is enabled for all dimensions in the input, even those that are not masked ($w_{i}(m_{1})\approx 1$). As a result, the variance of the noise in $\mathcal{M}_{2}$ scales as $\|w(m_{1})\|_{2}^{2}\leq d$. The more masking, the lower the variance. It may help to consider the change of variables $\sigma\leftarrow\sigma/\sqrt{d}$ in Equation 4, and $\sigma_{1,2}\leftarrow\sigma_{1,2}/\sqrt{d}$ in Proposition 2.4, to remove $d$ from $r^{\infty}_{X}$ and scale the noise variance with $d$. For RS (Equation 4), the noise variance scales as $d$. For ARS, only Equation 5 (the first step) suffers from variance scaled by $d$, while the second step’s variance (Equation 6) scales as $||w(m_{1})||_{2}^{2}$, which can be much smaller than $d$ when a large part of the image is masked. Reduced variance translates into higher accuracy, as well as $\underline{p_{+}}$ and $\overline{p_{-}}$ being further apart, for a larger $r^{\infty}_{X}$. 3. The variance reduction due to masking applies in the translation from the $B_{\infty}(r^{\infty})$ bound on the attack to the $B_{2}(r)$ sensitivity used in ARS. This variance reduction would not apply to an $L_{2}$-bounded adversary. Hence, our two-steps ARS architecture for $L_{\infty}$-bounded adversaries does not reduce to bounding $L_{\infty}$ with $L_{2}$ as the traditional RS application does, and our gains come explicitly from variance reduction enabled by adaptive masking.
@@ 4 Experiments
We evaluate on certified $L_{\infty}$ and standard test accuracy. Certified accuracy at radius $r^{\infty}$ is the percentage of test samples that are correctly classified and have an $L_{\infty}$ certificate radius $r^{\infty}_{X}\geq r^{\infty}$. Standard accuracy is obtained for $r^{\infty}=0$.
@@
Models We choose the standard ResNet (He et al., 2016) models as base classifiers $g$ with ResNet-110 for CIFAR-10 and ResNet-50 for CelebA and ImageNet. For ARS, our mask model $w$ is a simplified U-Net (Ronneberger et al., 2015) (see Appendix C for details). For the noise budget, we find that a fixed budget split performs reliably, and so in all experiments we split by $\sigma_{1}=\sigma_{2}=\sqrt{2}\sigma$.
Methods We compare to standard and strong static methods, design a baseline specifically for our masking approach, and evaluate the only sound input-dependent method prior to ARS. Cohen et al. is the standard approach to RS (Cohen et al., 2019). UniCR (Hong et al., 2022) learns the noise distribution for RS during training but is static during testing (while they propose an adaptive variant, it is not sound, so we restrict our comparison to the training variant). We tune hyper-parameters, and perform a grid search for $\beta$ (the parameter of the noise distribution) to maximize certified accuracy. We find $\beta=2.25$ (close to a Gaussian, with smaller tails) to perform best. Static Mask is our baseline that learns a fixed mask during training that does not adapt during testing. The mask is directly parameterized as pixel-wise weights that we multiply with the input and optimize jointly with the base classifier. Relative to ARS in Figure 1, this removes $\mathcal{M}_{1}$, sets $\sigma_{2}=\sigma$, and makes $w(.)=W$ static parameters rather than an adaptive prediction. Súkeník et al. (2021) conditions the variance $\sigma$ for RS on the input and is therefore test-time adaptive. We use code provided by the authors as is. Comparing ARS to the static baselines measures the value of test-time adaptivity, and comparing ARS to the variance adaptivity of Súkeník et al. measures the importance of more high-dimensional and expressive adaptation.
@@ 4.1 CIFAR-10 Benchmark: Classification with Distractor Backgrounds
Input dimension is a key challenge in $L_{\infty}$ certification using RS (see §2.4). We design our 20kBG benchmark to vary this parameter without affecting the task: we superimpose CIFAR-10 images onto a larger background from the $20k$ background images dataset Li et al. (2022a), split them into train and test sets, and resize to $k\times k$ pixels. The CIFAR-10 image is placed at random along the edges of the background image to maximize spatial variation. The spurious background increases the dimension of the input ($=k\times k\times 3$), making $L_{\infty}$ certification with RS more challenging, but it is uninformative for the task of CIFAR-10 by construction. Our mask model ($\mathcal{M}_{1}$) needs to learn to ignore the background to reduce the effective dimension of the input.
Table 1 summarizes the standard accuracy ($r=0$) of each approach on the different settings. In the top section, we vary $k$ (i.e. dimensionality) in $\{32,48,64\}$, where $k=32$ corresponds to original CIFAR-10 images. We keep $\sigma$ fixed to $0.75$ across these experiments. In next two rows, we vary $\sigma\in\{0.12,0.5,1.5\}$ at fixed $k=48$ and $k=64$, respectively. LABEL:fig:cifar-10-benchmark-plots shows the certified test accuracy at different radii $r^{\infty}$ for ARS and all baselines we consider.
@@
$\sigma=0.75$, $k=32$ 55.5 (0.8) 57.8 (0.8) 50.9 (5.9) 49.8 (1.0) 57.3 (0.6)
$\sigma=0.75$, $k=40$ 52.1 (0.4) 57 (0.7) 44.3 (1.2) 48.1 (1.6) 58.1 (1.8)
$\sigma=0.75$, $k=48$ 52.5 (1.2) 56 (0.8) 49.0 (2.2) 43.1 (3.1) 57.6 (1.5)
$\sigma=0.75$, $k=64$ 49.4 (2.0) 51.6 (1.0) - - 56.5 (2.6)
$\sigma=0.12$, $k=48$ 80.1 (0.8) 80.9 (0.7) 76.3 (1.9) 76.1 (4.1) 83.6 (0.4)
$\sigma=0.5$,  $k=48$ 64.3 (0.2) 64.1 (1.6) 54.0 (2.2) 55.9 (1.4) 66.0 (0.8)
$\sigma=1.5$,  $k=48$ 30.8 (0.2) 34.9 (2.4) 30.7 (0.9) 26.5 (0.7) 34.3 (0.2)
$\sigma=0.12$, $k=64$ 80.9 (1.0) 81.3 (1.0) - - 82.3 (1.0)
$\sigma=0.5$, $k=64$ 61.6 (2.7) 64 (1.4) - - 65.4 (1.8)
$\sigma=1.5$,  $k=64$ 28.3 (0.2) 25.9 (3.2) - - 30.6 (1.0)
Figure 10: ARS Masks on CIFAR-10 (20kBG) select the task-relevant input over the distractor background.
We make three observations. First, in most cases ARS outperforms all the baselines (static mask is slightly better with no distractor background, and with very high $\sigma$ when all methods have low accuracy) reaching a standard accuracy $4.9$ percentage points higher than the best baseline on our main setup ($k=64,\sigma=0.75$). Hong et al. (2022) and our own static mask baselines are the best performing, but all are close to Cohen et al. (2019) as these approaches provide little adaptivity. The improved standard accuracy of ARS translates to an improved certified accuracy at all certification levels (again except without distractor backgrounds, and with very high $\sigma$ when all methods have low accuracy). This is because ARS makes more accurate and confident predictions on more test examples, leading to a larger radius.
Second, as we grow the input dimension ($k$), the accuracy of ARS remains stable whereas that of baselines goes down, resulting in an increasing gap. For $k=32$ the best baseline (static mask) beats ARS by $0.5$ percentage points, whereas the gap between the same approaches reaches $1.6$ percentage points in favor of ARS when $k=48$, and $4.9$ points at $k=64$. This is because ARS’ mask is able to rule out spurious background information, reducing the noise in the second step, as shown on Figure 10. Thanks to this masking, ARS is much less sensitive to increases in dimensionality (see Table 1).
Third, we observe that ARS does much better than the baselines in low to moderate noise regimes ($\sigma=0.5,0.75$) whereas it performs similarly or even slightly worse in high noise regimes ($\sigma=1.5$). This is because in the high noise regime, the first ARS step is too noisy for the mask to be efficient.

### v3

# Adaptive Randomized Smoothing: Certified Adversarial Robustness for Multi-Step Defences

@@ Abstract
We propose Adaptive Randomized Smoothing (ARS) to certify the predictions of our test-time adaptive models against adversarial examples. ARS extends the analysis of randomized smoothing using $f$-Differential Privacy to certify the adaptive composition of multiple steps. For the first time, our theory covers the sound adaptive composition of general and high-dimensional functions of noisy inputs. We instantiate ARS on deep image classification to certify predictions against adversarial examples of bounded $L_{\infty}$ norm. In the $L_{\infty}$ threat model, ARS enables flexible adaptation through high-dimensional input-dependent masking. We design adaptivity benchmarks, based on CIFAR-10 and CelebA, and show that ARS improves standard test accuracy by $1$ to $15\%$ points. On ImageNet, ARS improves certified test accuracy by up to $1.6\%$ points over standard RS without adaptivity. Our code is available at https://github.com/ubc-systopia/adaptive-randomized-smoothing.
@@ 1 Introduction
Despite impressive accuracy, deep learning models still show a worrying susceptibility to adversarial attacks. Such attacks have been shown for a large number of tasks and models (Costa et al., 2023; Chakraborty et al., 2018), including areas where security and safety are critical such as fraud detection (Pumsirirat and Liu, 2018) or autonomous driving (Cao et al., 2021).
@@
We evaluate our adaptive randomized smoothing method in three settings (§4). For image classification, we first design a challenging benchmark based on CIFAR-10, and we show that ARS can improve accuracy by up to $15\%$. For spatially-localized face attribute classification on the CelebA dataset, we show that ARS improves accuracy by up to $7.7\%$. For large-scale image classification on ImageNet, ARS maintains accuracy and improves certified accuracy by up to $1.6\%$.
@@ 2.1 Related Work: Adversarial Robustness, Randomized Smoothing, and Differential Privacy
While sound, RS is static during testing even though attacks may adapt. Recent work aims to make RS adapt at test time (Súkeník et al., 2021; Alfarra et al., 2022a; Hong et al., 2022). While pioneering, these works are restricted in either their soundness or their degree of adaptation and resulting improvement. Súkeník et al. (2021) soundly adapt the variance for RS by the distance between test and train inputs. However, this only provides minimal adaptivity, with only minor improvement to certification. Alfarra et al. (2022a) adapt the variance for RS to each test input, but the analysis is not end-to-end, and hence not sound (Súkeník et al., 2021) (except for their memory-based edition, but this requires storage of all examples). UniCR (Hong et al., 2022) adapts the noise distribution for RS, primarily to the data distribution during training, and optionally to input during testing. The train-time adaptation is sound, but the test-time adaptation is not due to the same issue raised by Súkeník et al. (2021). We propose ARS to advance certified test-time adaptivity: our approach is sound and high-dimensional to flexibly adapt the computation of later steps conditioned on earlier steps by differential privacy theory.
@@
$\boldsymbol{f}$-DP (Dong et al., 2019) is a notion of privacy that extends $(\epsilon,\delta)$-DP, and defines privacy as a bound on the power of hypothesis tests. Appendix A provides more details on $f$-DP. The main result we leverage is Theorem 2.7 of Dong et al. (2022) that for a Gaussian mechanism $\mathcal{M}(X)=\theta(X)+z$, $z\sim\mathcal{N}(0,\frac{r^{2}}{\mu^{2}})$, such that for any neighbouring $X,X^{\prime}$, $\theta(X)-\theta(X^{\prime})\in B_{2}(r)$ (i.e., the $L_{2}$ sensitivity of $\theta$ is $r$), we have that $\mathcal{M}$ is $G_{\mu}$-DP with the function $G_{\mu}$ defined as:
$G_{\mu}(\alpha)=\Phi\Big{(}\Phi^{-1}(1-\alpha)-\mu\Big{)},\textrm{ for }\alpha\in[0,1].$ (2)
We leverage two key properties of $f$-DP. First, $f$-DP is resilient to post-processing. That is, if mechanism $\mathcal{M}$ is $f$-DP, $\textrm{proc}\circ\mathcal{M}$ is also $f$-DP. Second, $f$-DP is closed under adaptive composition. We refer to §3 in Dong et al. (2022) for the precise definition and use their Corollary 3.3: the adaptive composition of two Gaussian mechanisms $G_{\mu_{1}}$-DP and $G_{\mu_{2}}$-DP is itself $G_{\mu}$-DP with
$\mu=\sqrt{\mu_{1}^{2}+\mu_{2}^{2}}.$ (3)
@@ 2.4 ARS against $L_{\infty}$-Bounded Adversaries
Important remarks. 1. $w(.)$ is a masking function, adaptively reducing (if $w_{i}(m_{1})\ll 1$) the value of $X_{i}$ and thereby the attack surface of an $L_{\infty}$ attacker. This reduces the effective dimension of the input to $\mathcal{M}_{2}$. 2. Reducing the dimension enables a reduction in the noise variance in $\mathcal{M}_{2}$, at fixed privacy guarantee $G_{\mu_{2}}$. The variance reduction is enabled for all dimensions in the input, even those that are not masked ($w_{i}(m_{1})\approx 1$). As a result, the variance of the noise in $\mathcal{M}_{2}$ scales as $\|w(m_{1})\|_{2}^{2}\leq d$. The more masking, the lower the variance. It may help to consider the change of variables $\sigma\leftarrow\sigma/\sqrt{d}$ in Equation 4, and $\sigma_{1,2}\leftarrow\sigma_{1,2}/\sqrt{d}$ in Proposition 2.4, to remove $d$ from $r^{\infty}_{X}$ and scale the noise variance with $d$. For RS (Equation 4), the noise variance scales as $d$. For ARS, only Equation 5 (the first step) suffers from variance scaled by $d$, while the second step’s variance (Equation 6) scales as $||w(m_{1})||_{2}^{2}$, which can be much smaller than $d$ when a large part of the image is masked. Reduced variance translates into higher accuracy, as well as $\underline{p_{+}}$ and $\overline{p_{-}}$ being further apart, for a larger $r^{\infty}_{X}$. 3. The variance reduction due to masking applies in the translation from the $B_{\infty}(r^{\infty})$ bound on the attack to the $B_{2}(r)$ sensitivity used in ARS. This variance reduction would not apply to an $L_{2}$-bounded adversary (an attack that only changes pixels with mask values of one yields no sensitivity improvement). Hence, our two-steps ARS architecture for $L_{\infty}$-bounded adversaries does not reduce to bounding $L_{\infty}$ with $L_{2}$ as the traditional RS application does, and our gains come explicitly from variance reduction enabled by adaptive masking against an $L_{\infty}$ attack.
@@ 4 Experiments
We evaluate on standard and $L_{\infty}$-certified test accuracy. Certified accuracy at radius $r^{\infty}$ is the percentage of test samples that are correctly classified and have an $L_{\infty}$ certificate radius $r^{\infty}_{X}\geq r^{\infty}$. Standard accuracy is obtained for $r^{\infty}=0$.
@@
Models We choose the standard ResNet (He et al., 2016) models as base classifiers $g$ with ResNet-110 for CIFAR-10 and ResNet-50 for CelebA and ImageNet. For ARS, our mask model $w$ is a simplified U-Net (Ronneberger et al., 2015) (see Section C.1 for details). For the noise budget, we find that a fixed budget split performs reliably, and so in all experiments we split by $\sigma_{1}=\sigma_{2}=\sqrt{2}\sigma$.
Methods We compare to standard and strong static methods, design a baseline specifically for our masking approach, and evaluate the only sound input-dependent methods prior to ARS. Cohen et al. is the standard approach to RS (Cohen et al., 2019). UniCR (Hong et al., 2022) learns the noise distribution for RS during training but is static during testing (while they propose an input-adaptive variant, it is not sound so we restrict our comparison to the training variant). We tune hyper-parameters, and perform a grid search for $\beta$ (the parameter of the noise distribution) to maximize certified accuracy. We find that $\beta=2$ (Gaussian), or $\beta=2.25$ (close to a Gaussian, with smaller tails), to perform best in high and low $(k,\sigma)$ values, respectively (see Appendix D) for details. Static Mask is our baseline that learns a fixed mask during training that does not adapt during testing. The mask is directly parameterized as pixel-wise weights that we multiply with the input and optimize jointly with the base classifier. Relative to ARS in Figure 1, this removes $\mathcal{M}_{1}$, sets $\sigma_{2}=\sigma$, and makes $w(.)=W$ static parameters rather than an adaptive prediction. Súkeník et al. (2021) conditions the variance $\sigma$ for RS on the input and is therefore test-time adaptive. We use code provided by the authors as is. Comparing ARS to the static baselines measures the value of test-time adaptivity, and comparing ARS to the variance adaptivity of Súkeník et al. measures the importance of more high-dimensional and expressive adaptation.
@@ 4.1 CIFAR-10 Benchmark: Classification with Distractor Backgrounds
Input dimension is a key challenge in $L_{\infty}$ certification using RS (see §2.4). We design our 20kBG benchmark to vary this parameter without affecting the task: we superimpose CIFAR-10 images onto a larger distractor background from the $20k$ background images dataset Li et al. (2022a). The backgrounds are split into train and test sets, and resized to $k\times k$ pixels (where $k\geq 32$). The CIFAR-10 image (of dimensions $32\times 32\times 3$) is then placed at random along the edges of the background image to maximize spatial variation. The spurious background increases the dimension ($=k\times k\times 3$) of the input when $k>32$, making $L_{\infty}$ certification with RS more challenging, but is uninformative for the task of CIFAR-10 by construction. Our mask model ($\mathcal{M}_{1}$) needs to learn to ignore the background to reduce the effective dimension of the input. For computational reasons, we run all certification results on a 200 sample subset of the CIFAR-10 test set. Appendix D shows extended results, details about hyperparameter tuning (C.3), and results on larger test-sets (D.1). We set the failure probability of the certification procedure to 0.05, use 100 samples to select the most probable class, and 50,000 samples for the Monte Carlo estimate of $\underline{p_{+}}$.
Table 1 summarizes the standard accuracy ($r=0$) of each approach on the different settings. We vary $\sigma$ in $\{0.25,0.5,1.0\}$ and $k$ in $\{32,48,64,96\}$ to show the effect of noise levels and dimensionality on the accuracy of different approaches ($k=32$ corresponds to original CIFAR-10 images). LABEL:fig:cifar-10-benchmark-plots shows the certified test accuracy at different radii $r^{\infty}$ for ARS and all baselines we consider.
@@
$\sigma=0.25,\ k=32$ 70.6 (1.1) 73.9 (0.8) 69.8 (1.4) 68.6 (2.8) 72.6 (0.9)
$\sigma=0.5,\ \ \ k=32$ 63.6 (2.0) 64.8(0.9) 62.8 (1.4) 59.1 (1.6) 64 (1.4)
$\sigma=1.0,\ \ \ k=32$ 48 (0.7) 47.3 (1.3) 46.1 (0.9) 44.6 (1.0) 49.3 (0.6)
$\sigma=0.25,\ k=48$ 71.6 (1.0) 72 (2.0) 69.7 (0.8) 65 (0.7) 75.5 (1.0)
$\sigma=0.5,\ \ \ k=48$ 64.3 (0.2) 64.1 (1.6) 60.3 (0.6) 53.5 (1.8) 66 (1.6)
$\sigma=1.0,\ \ \ k=48$ 42.5 (2.1) 45.1 (1.2) 44.3 (0.2) 34.1 (1.0) 47.6 (2.0)
$\sigma=0.25,\ k=64$ 71.6 (0.9) 73.1 (3.2) 67.8 (0.5) 64.1 (0.8) 77 (1.8)
$\sigma=0.5,\ \ \ k=64$ 63 (1.6) 62.5 (1.7) 58.7 (0.2) 45.1 (1.1) 69.9 (1.2)
$\sigma=1.0,\ \ \ k=64$ 41.3 (1.8) 40.0 (0.5) 42.2 (0.6) 26.5 (0.7) 50.4 (2.5)
$\sigma=0.25,\ k=96$ 65.3 (1.6) 71.8 (1.3) 68.8 (1.8) 45.5 (0.9) 78.3 (2.2)
$\sigma=0.5,\ \ \ k=96$ 56.6 (2.4) 59.5 (1.4) 59.7 (1.3) 10.8 (2.3) 69.8 (1.2)
$\sigma=1.0,\ \ \ k=96$ 33.8 (3.8) 36.9 (0.5) 41.3 (2.4) 10.4 (0.3) 56.3 (2.3)
Table 1: Standard Accuracy ($r=0$) on CIFAR-10 (20kBG). Our 20kBG benchmark places CIFAR-10 images on larger background images. We report the mean accuracy and standard deviation over three seeds. ARS achieves higher accuracy across noise $\sigma$ and input dimension $k$ (Δ indicates adaptivity). We provide results with more $\sigma$ levels in Appendix D.
$k$ Input Images ARS Masks
32 Refer to caption Refer to caption
64 Refer to caption Refer to caption
96 Refer to caption Refer to caption
Figure 7: (left) Original CIFAR-10 images superimposed on backgrounds for different $k$ (except $k=32$ which is no background), and (right) their corresponding masks (grayscale) inferred by our mask model $w$. All masks are for $\sigma=0.5$. Section D.2 shows all the corresponding images across our multi-step architecture.
We make three observations. First, ARS outperforms all the baselines under distractor backgrounds ($k>32$). Static mask is slightly better at $k=32$, probably because CIFAR-10 images lack enough "irrelevant" information for ARS to discard. This is precisely why we introduce this benchmark, where we explicitly add such irrelevant information to input images. Indeed, at $k>32$, we observe that the standard accuracy of ARS improves, translating to an improved certified accuracy at all certification levels, since more accurate and confident predictions give a larger certified radius $r^{\infty}$.
Second, as we grow the input dimension $k$, the accuracy of ARS remains either stable or increases, whereas that of other baselines goes down, resulting in an increasing gap. For instance, at $\sigma=1.0$, the gap between ARS and the best baseline is $1.3$% points for $k=32$, $2.5$% points for $k=48$, $8.2$% points for $k=64$ and $15$% points for $k=96$. As $k$ grows, the amount of relevant information (a $32\times 32\times 3$ CIFAR-10 image) remains the same, whereas the amount of spurious background information increases. ARS’ mask is able to rule out spurious pixels, reducing the noise in the second step (Figure 7). Thanks to this masking, ARS is much less sensitive to increases in dimensionality.
Third, we observe that except $k=32$, ARS improves over all the baselines in low ($\sigma=0.25$) to high ($\sigma=1.0$) noise regimes. In fact, this trend continues to persist in higher noise regimes (Appendix D). Similar to previous observations, we notice that as we increase $\sigma$, other baselines’s accuracy drops significantly whereas ARS accuracy drops much less, displaying higher noise tolerance.